OpenStudy (anonymous):

A line through which two points would be perpendicular to the line y = 2/3x - 5/2? A. (–4, 3) and (–2, –6) B. (–4, –3) and (2, 6) C. (–4, 3) and (2, –6) D. (4, –3) and (2, 6)

OpenStudy (anonymous):

\[y = \frac{ 2 }{ 3 }x - \frac{ 5 }{ 2 }\]

OpenStudy (texaschic101):

can you tell me the slope of y = 2/3x - 5/2 ?

OpenStudy (texaschic101):

In y = mx + b form, the slope will be the number in the m position. Can you tell me the slope now ?

OpenStudy (anonymous):

 My drawing sucks but 3/1 I think